OpenStudy (anonymous):

if 2f(x) - 3f(1/x) = x^2, determine f(2) (show steps please)

OpenStudy (anonymous):

just plug in 2 for x.

OpenStudy (anonymous):

its not that simple

OpenStudy (anonymous):

plug in 2, then plug in 1/2, and you will have a system of equations in two unknowns, one will be f(2), the other will be f(1/2), and you can solve for f(2).

OpenStudy (anonymous):

\[2f(2)-3f(\frac{1}{2})=4, -3f(2)+2f(\frac{1}{2})=\frac{1}{4} \] now we have to solve this system.

OpenStudy (anonymous):

You end up with f(2)=-7/4